Suppose that the distance of fly balls hit to the outfield (in baseball) is normally distributed with a mean of 250 feet and a standard deviation of 51 feet.
Use your graphing calculator to answer the following questions. Write your answers in percent form. Round your answers to the nearest tenth of a percent.
a) If one fly ball is randomly chosen from this distribution, what is the probability that this ball traveled fewer than 198 feet?PP

  1. (fewer than 198 feet) = %

b) If one fly ball is randomly chosen from this distribution, what is the probability that this ball traveled more than 210 feet?PP

  1. (more than 210 feet) =
Answer #1

a) P(Fewer than 198 feet)

= P(z < (198 - 250)/51)

= P(z < -1.02)

= 15.4 %

b) P(More than 210 feet)

= P(z > (210 - 250)/51)

= P(z > -0.78)

= 78.4 %